OpenStudy (anonymous):

A rubber balloon has become negatively charged, and the charge is measured as 1.00 * 10^-14C according to the this charge, the balloon has an excess of how many electrons?

OpenStudy (akashdeepdeb):

Charge on balloon = \(- 10^{-14}C\) Thus, as it is negatively charged, we know that it has an excess of electrons. 1 electron charge = \(-1.6 * 10^{-19} C\) Thus if one electron has \(-1.6 * 10^{-19} C\) charge. Then let x electrons have \(- 10^{-14}C\) charge. Therefore, more electrons means more charge thus they have a direct relation. Thus,\[\frac{x}{1} = \frac{- 10^{-14}C}{-1.6 * 10^{-19} C}\] \[x = 62500\] And therefore we can conclude, that the number of extra electrons on the ballon = 62500\(e\) Understood? :)

OpenStudy (anonymous):

thank you so much

OpenStudy (akashdeepdeb):

:)